OpenStudy (anonymous):

If 20 lb of rice and 10 lb of potatoes cost $16.20 and 30 lb of rice and 12 lb of potatoes cost $23.04, how much will 10 lb of rice and 50 lbs of potatoes cost? Let x = price of rice per pound Let y = price of potaotes per pound 20x + 10y = 16.2 30x + 12y = 23.04 Solve the system of equations for x and y. Then substitute x and y in 10x + 50y and evaluate the expression. 20x+10y=16.2 R1 30x+50y=23.04 R2

OpenStudy (precal):

Let x = price of rice per pound Let y = price of potaotes per pound 20x + 10y = 16.2 30x + 12y = 23.04 this is correct just solve for x and y

OpenStudy (anonymous):

@precal thats where i am having problems.. i am not sure on how to solve for x and y.. the book has me confused with rows

OpenStudy (phi):

the way to solve 2 equations with 2 unknowns is to "eliminate" one of the variables 20x+10y=16.2 30x+50y=23.04 Let's try to get rid of the y's if the first equation had -50y (instead of 10y), when you added the 2 equations together, the -50y + 50y become 0 (and y is gone). Notice that you can multiply both sides of the 1st equation by any number you want, and it will remain true. what number should you multiply it by to make the 10y become -50y ?

OpenStudy (anonymous):

-5

OpenStudy (phi):

yes, so multiply the 1st equation by -5 what do you get?

OpenStudy (anonymous):

-5(20x+10y)=16.2 distribute -5 to both the 20 and the 10? -50x-50y=16.2

OpenStudy (phi):

yes, distribute the -5 and you have to multiply both sides if we had (1+2)= 3 (which is true) and we multiply by -5 -5(1+2) = -5*3 -5*1 + -5*2 = -15 -5 -10 = =15 -15= -15 it is still true. but if we do what you did it would not work

OpenStudy (anonymous):

thats not right..

OpenStudy (anonymous):

what i had i mean

OpenStudy (phi):

yes I know. start again: multiply both sides of the = by -5 when you distribute the -5 that means multiply *everything* inside the parens not just the y)

OpenStudy (anonymous):

-100x-50y=-81

OpenStudy (phi):

looks good. we now have -100x-50y=-81 30x+50y=23.04 add the 2 equations

OpenStudy (anonymous):

-70x+0y=-57.96 but that doesnt look right to me

OpenStudy (phi):

0y is just 0 so you have -70x = -57.96 divide both sides by -70

OpenStudy (anonymous):

.828

OpenStudy (phi):

Well, that was a nice exercise, but after you mentioned it didn't look right I re-read the question. the two equations we should start with are 20x + 10y = 16.2 30x + 12y = 23.04

OpenStudy (phi):

Notice there isn't a nice number to multiply the first equation to get a -12y (but there is a number, it is -12/10 or -1.2) try that. multiply the first equation by -1.2

OpenStudy (anonymous):

ok. -24x+-12y=-19.44

OpenStudy (phi):

now add the two equations -24x+-12y=-19.44 30x + 12y = 23.04

OpenStudy (anonymous):

i got .6 or 3/5

OpenStudy (phi):

x = price of rice per pound so I would call it $0.60 (60 cents) now use one of the 2 equations to find y replace x with 0.6 and solve for y It does not matter which, so use this one 20x+10y=16.2

OpenStudy (anonymous):

.42

OpenStudy (phi):

Finally, how much will 10 lb of rice and 50 lbs of potatoes cost?

OpenStudy (anonymous):

27 dollars

OpenStudy (phi):

how did you get that?

OpenStudy (anonymous):

i took my 10 pounds of rice and multiplied it by 60 cents and the 50 lbs of potatoes and multiplied by .42 then added the 2 together

OpenStudy (phi):

yes, excellent. Now here is a question. How did I pick -1.2 to multiply the first equation? 20x + 10y = 16.2 30x + 12y = 23.04 I saw the 10y and said, if that were 1y I could multiply by -12 and get -12y and -12y plus 12y in the other equation will add to 0. then I said, if I divide by 10 then 10y will become 1y so, putting those together I get -12/10 or -1.2 for the multiplier. Can you pick a multiplier to get rid of the x's instead of the y's ?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

well a neg 3

OpenStudy (phi):

-3 *20x = -60x and that does not exactly cancel out +30x in the other equation. read my reasoning carefully.

OpenStudy (anonymous):

-1.5?

OpenStudy (phi):

yes. -1.5*20x= -30 x and added to +30x in the other equation gives 0! I would get -1.5 by saying : divide 20x by 20 to get 1x. multiply 1x by -30 to get -30x both steps together are : -30/20 or -1.5

OpenStudy (phi):

Now you can solve 2 equations and 2 unknowns.

OpenStudy (anonymous):

and that works for every problem?

OpenStudy (anonymous):

i think you should rewrite my math book and have them publish it instead of what i have.. because you make so much more sense

OpenStudy (phi):

yes, unless it's not solvable to begin with. (sometimes both equations are the same equation in disguise and that is a problem) or sometimes the equations don't have a solution. But if there is an answer, this way always works.

OpenStudy (phi):

thanks.

OpenStudy (anonymous):

is there an easy way to tell if it is not solvable

OpenStudy (phi):

yes, sometimes when you eliminate x you will also eliminate y at the same time and you get 0+0 = right side if the right side is not 0, there is no solution. 0= number is never true. if the right side is 0, then that means the 2 equations are really the same equation and the solution is any x,y pair on the line (lots of solutions) but in general, the two equations (there are lines) cross somewheres, and there is only one (x,y) pair that is on both lines (satisfy both equations)

OpenStudy (anonymous):

and if 0=0 ?

OpenStudy (anonymous):

how can u do this problem on a graphing calc?

OpenStudy (phi):

and if 0=0 ?the solution is any x,y pair on the line (lots of solutions) how can u do this problem on a graphing calc? both equations are lines, and you graph them and see where they meet. Here is your problem http://www.wolframalpha.com/input/?i=20x+%2B+10y+%3D+16.2%2C++30x+%2B+12y+%3D+23.04

OpenStudy (anonymous):

i have never heard of that website.. that helps alot

OpenStudy (precal):

if you graph your equations, you will get 1 of 3 cases: Case 1: lines will cross in one place, that is known as one solution Case 2: lines are parallel and therefore do not cross (known as no solution) Case 3: lines are the same, therefore cross in many places (known as many solutions) hope that helps